# Galactic Archaeology with Gaia

### v1

@@ 3.4 Globular clusters
The premise of the study by Kruijssen et al. (2019b) is identical to the starting point in Massari et al. (2019), namely that in-situ and accreted populations behave differently in the age-metallicity space. Instead of linking the chrono-chemical properties of the Milky Way globulars to their orbital behaviour, Kruijssen et al. (2019b) attempt to decipher the make-up of the Galactic GCs using the cluster E-MOSAICS models based on the EAGLE Cosmological numerical simulations (Pfeffer et al., 2018; Kruijssen et al., 2019a). Using the E-MOSAICS framework, Kruijssen et al. (2019b) isolate the 3 most significant events in the Milky Way’s past, contributing the largest numbers of accreted GCs. However, without the use of the orbital information, identifying the progenitors of these systems turns out to be rather difficult. While the smallest group associated with the Sgr dwarf appears to be self-evident, the two most massive events are much more uncertain. Some 22 GCs are attributed to i) the so-called Canis Major dwarf (see Martin et al., 2004) and ii) the hypothetical Kraken galaxy, predicted by Kruijssen et al. (2019b) to be the largest event in the Milky Way’s life, bringing with it, some $\sim 2\times 10^{9}M_{\odot}$ of stellar mass. The current consensus on the accreted nature of the Canis Major is rather skeptical given the disc origin of the various “streams” associated with it (see e.g. Xu et al., 2015; Price-Whelan et al., 2015; de Boer et al., 2018; Sheffield et al., 2018; Laporte et al., 2020). Most importantly, however, neither of the two groups of Galactic GCs identified by Kruijssen et al. (2019b) exhibits coherent orbital behavior and thus both are instead most likely amalgamations of parts of different mergers. The analysis of Kruijssen et al. (2019b) is trail-blazing but falls short due to the limited quality of the age-metallicity information for the Galactic GCs and the neglect of the orbital information. An attempt to rectify the latter is made in Kruijssen et al. (2020), which re-uses the IOM classification made by Massari et al. (2019). In a similar vein, Forbes and Bridges (2010) starts with GC groups constructed by Massari et al. (2019) and tweaks them very slightly using the age-metallicity information at hand.
Figure 14: Groups of GCs in action-angle space inferred from chemodynamical modeling. The coloured symbols are the observed GCs: different colours indicate different groups, with each GC coloured by its most likely group. The accreted groups of GCs are indicated with their $1-,2-,3-\sigma$ confidence intervals. [Reproduced from Callingham et al. (2022)].
@@
Compared to all preceding attempts, the study of Callingham et al. (2022) stands out as the only unbiased GC classification scheme. Instead of hand-picking groups of clusters in the chosen parameter space, they devise a method similar to Gaussian Mixture Modeling. For the Milky Way GCs, ages, metallicities, energy, and three orbital actions are modeled with component mixtures to obtain objective membership probabilities for each GC (see Fig. 14). This innovative approach is demonstrated to work well for simulated data. Unfortunately, Callingham et al. (2022) also reports a large degree of overlap between GCs across all dimensions studied. To mitigate against the arising degeneracies, Callingham et al. (2022) decide to initialize their mixture models to the GC classes reported by Massari et al. (2019).
Concerned with the separation between the GCs born in-situ and those formed in dwarfs and later accreted, Belokurov and Kravtsov (2023b) go back to the idea of the “critical energy” highlighted in Myeong et al. (2018c). They show that in the stellar halo too, a rather sharp boundary exists in $E,L_{z}$ space which divides stars into mostly accreted and mostly in-situ. This division is mapped by the clear change in the overall levels of [Al/Fe] ratio amongst the filed stars. Using the $E,L_{z}$ boundary delineated by the [Al/Fe] levels, Belokurov and Kravtsov (2023a) classify all Milky Way GCs into two classes and find 106 out of 164 to be of in-situ origin, and 58 to be accreted. Therefore the most up-to-date GS/E membership of $\sim 20$ GCs translates into $\sim 1/3$ of the total accreted cluster population in the Galaxy.
@@ 3.5 Simulations
While the above reflects a more tailored approach to the simulations of a Milky Way-GS/E merger, other works have focused on large suites of Milky Way-mass galaxies simulated in a cosmological context. In this case, an ‘ideal’ Milky Way-GS/E system is harder to find, but such analyses are essential in order to put the Milky Way assembly history in context with the wider galaxy population. Fattahi et al. (2019) use the Auriga suite of simulations to show that approximately one-third of the (28) haloes have a dominant stellar halo component similar to the observed Milky Way. From this subset, Fattahi et al. (2019) deduce that these merger events typically occur 6-10 Gyr ago, with merging fragments of stellar mass $M_{\rm star}\sim 10^{9}-10^{10}M_{\odot}$. Orkney et al. (2023) extend the analysis of the GS/E events in Auriga and highlight the diversity of possible progenitor systems, with a broad range of stellar, dark matter masses and rotational properties. Mackereth et al. (2019) use the EAGLE simulation suite to show that merger events with a similar pattern in chemistry and dynamics to the real GS/E are typically related to massive dwarfs with $M_{\rm star}\sim 10^{8.5}-10^{9}M_{\odot}$. Moreover, they argue that the high eccentricity of the GS/E stars today suggests the merger would have taken place at $z\lesssim 1.5$. One particular Milky Way analog within the EAGLE suite is additionally studied in greater detail by Bignone et al. (2019). Elias et al. (2020) show that many of the conclusions reached through studies of GS/E-like events in suites of zoom-in simulations also hold in a much larger but lower-resolution sample from Illustris. Finally, Dillamore et al. (2022) uses the ARTEMIS simulations to again analyse GS/E-like events in cosmological simulations. In agreement with Fattahi et al. (2019) they find that typically a third of the Milky Way-mass systems contain a feature similar to the GS/E.
@@
Moving forward, a combined effort of tailored Milky Way-GS/E mergers (e.g. Koppelman et al., 2020; Naidu et al., 2021) and analyses of cosmological simulation suites (e.g. Fattahi et al., 2019; Mackereth et al., 2019; Dillamore et al., 2022) are needed to (a) link the observed properties of the GS/E to its initial state and co-evolution with the Milky Way, and (b) calibrate our various observational techniques to infer these properties. An important step in the right direction are the works of Amarante et al. (2022) and Rey et al. (2023) who create hydro-dynamical models of the GS/E-Milky Way encounter. Amarante et al. (2022) build tailor-made single merger models with realistic star-formation recipes and detailed chemistry. They show that many previously detected halo substructures can indeed be attributed to the single GS/E-like event. Rey et al. (2023) pioneers a new powerful and flexible way to carry out inference with Cosmological hydro-dynamical simulations. They use genetic modifications of the simulation’s initial conditions (see Roth et al., 2016; Rey and Pontzen, 2018) to modify the properties (such as the mass ratio) of the GS/E merger. Rey et al. (2023) take a Milky Way zoom-in simulation from the VINTERGATAN suite and show how the properties of both the host and the resulting debris cloud change as the merger strength is varied. The Milky Way suffers more dramatic truncation and becomes more bulge-dominated as the GS/E mass is increased (see Fig. 15). In the absence of detailed chemistry, Rey et al. (2023) find a pronounced kinematic degeneracy for the stellar halo: different mergers and different mixtures of accreted and in-situ stars end up producing similar looking radially-biased “sausage”-like structures.
@@ 3.6.2 The mass budget of the GS/E progenitor
The most recent stellar mass estimates of the GS/E are typically a few ×$\times 10^{8}M_{\odot}$ (e.g. Naidu et al., 2021; Lane et al., 2023), which is almost an order of magnitude lower than earlier estimates of $\sim 10^{9}M_{\odot}$ (e.g. Belokurov et al., 2018b; Helmi et al., 2018; Feuillet et al., 2020). This is likely due to a greater purity of GS/E stars in more recent studies, but, even now, a robust stellar mass estimate is lacking. Nonetheless, a lower mass estimate for GS/E has several important consequences. For example, it would imply less damage to the pre-existing disc — this would agree with observations that the high-$\alpha$ disc is still intact. However, simulated analogues of the Milky Way-GS/E encounter would need to be updated in light of this seemingly preferred lower GS/E mass.
@@
Finally, it is also worth considering what the gas content of the GS/E progenitor was. The early nature of the merger strongly suggests that the progenitor was gas-rich, but where and when was the gas lost? Could it contribute to (or re-start) the Galactic star formation? Does the two infall model of the Galaxy (Chiappini et al., 1997) still hold weight in light of the apparent gas-rich GS/E merger?
@@ 3.6.3 Initial state of the progenitor
Although much work is needed to characterize the GS/E debris today, ultimately we also want to know what the initial state of the progenitor was (i.e. before any interactions with the Milky Way). How much information is encoded in the $z=0$ debris? For example, if the progenitor was a disc galaxy with a large satellite (Sequoia?), broadly similar to the LMC and SMC, it could produce a variety of halo substructures with distinct properties. However, even without a complex morphology and kinematics, high-mass events tend to break up into a large number of clumps of the $E,L_{z}$ space (see e.g. Jean-Baptiste et al., 2017; Belokurov et al., 2023). Indeed, the ‘chevrons’ in $r-V_{r}$ space uncovered by Belokurov et al. (2023) could result from the phase-mixing of the massive GS/E. However, Dillamore et al. (2023) recently showed that such stellar substructures in the local halo could also generated by bar resonances.
@@ 3.6.4 Debris distribution in the Galactic halo
Most of the analysis of the GS/E debris has been limited to the inner ($r\lesssim 30$ kpc) halo. This, in part, is because here we have the richest multi-dimensional data, but also because the high eccentricity of the GS/E orbit implies there are plenty of GS/E stars at small Galactic radii. However, there is not much evidence (so far) for GS/E stars at low total $E$. How deep a satellite can sink in the halo depends not only on its mass, but also on (i) the density of the host and the satellite, and (ii) the initial amount of angular momentum (Amorisco, 2017; Vasiliev et al., 2022). The low $E$ regime is complicated by the overwhelming in-situ population, but there is potentially a rich amount of uncharted information about the GS/E lurking in the Milky Way’s depths.
At the other ‘end’ of the Galaxy, there is still much to learn about the GS/E stars at large distances. Here, shells and earlier stripped GS/E stars are likely prevalent in the halo. Indeed, the distant ‘echoes’ of the GS/E are now starting to be discovered (Chandra et al., 2023a), and with increasing amounts of data at large Galactic radii, this will likely be a fruitful expedition in future studies. Mapping the GS/E debris over a wide range of radii (or energies) will be vital in order to reconstruct the full dynamical history of the Milky Way-GS/E interaction.
@@ 3.6.5 A complete re-interpretation
Although the vast majority of the works described above advocate (or assume?) that the GS/E is a single, early, and massive progenitor, this is not the only interpretation in the literature. In particular, in a series of recent papers (Donlon et al., 2022; Donlon and Newberg, 2023; Donlon et al., 2023), Donlon et al. argue that the GS/E in fact comprises debris from several lower mass radial mergers, which moreover, were accreted fairly recently. This alternative view, coupled with the large uncertainty in several of the GS/E properties (e.g. total stellar mass, time of accretion) shows that we are still far from a robust characterization of this stellar halo component. Perhaps the best way to avoid the degeneracy between total mass and the number of progenitors is the use of the well-known stellar mass-metallicity (MZR) relation (as several lower mass fragments will have lower metallicity than a single higher mass system). However, the biases associated with different selection methods (see e.g. Carrillo et al., 2024) coupled with uncertainties in the time evolution of the MZR has hampered this crucial line of evidence (see also Section 5.3). Nonetheless, it is clear that a scenario accounting for all of the chemodynamical evidence is required in order to fully solve the GS/E puzzle.
@@ 4.1 In-situ stellar halo
Figure 16: Left: Azimuthal velocity ($V_{\phi}$) as a function of age for local metal-rich stars selected from Gaia DR2. Middle: Distribution of ages for stars in different bins of $V_{\phi}$ and [Fe/H]. Disc stars (red) are metal-rich with high $V_{\phi}$, Splash (black) is metal-rich with low $V_{\phi}$, and GS/E (blue) stars are metal-poor(er) with low $V_{\phi}$. Right: Cumulative age distribution. Belokurov et al. (2020) argue that the truncated age distribution of the Splash stars (indicated by a vertical thick red line) can put a strong constraint on the time of the last major merger. [Adapted from Belokurov et al. 2020].

### v2

@@ 3.4 Globular clusters
The premise of the study by Kruijssen et al. (2019b) is identical to the starting point in Massari et al. (2019), namely that in-situ and accreted populations behave differently in the age-metallicity space. Instead of linking the chrono-chemical properties of the Milky Way globulars to their orbital behaviour, Kruijssen et al. (2019b) attempt to decipher the make-up of the Galactic GCs using the cluster E-MOSAICS models based on the EAGLE Cosmological numerical simulations (Pfeffer et al., 2018; Kruijssen et al., 2019a). Using the E-MOSAICS framework, Kruijssen et al. (2019b) isolate the 3 most significant events in the Milky Way’s past, contributing the largest numbers of accreted GCs. However, without the use of the orbital information, identifying the progenitors of these systems turns out to be rather difficult. While the smallest group associated with the Sgr dwarf appears to be self-evident, the two most massive events are much more uncertain. Some 22 GCs are attributed to i) the so-called Canis Major dwarf (see Martin et al., 2004) and ii) the hypothetical Kraken galaxy, predicted by Kruijssen et al. (2019b) to be the largest event in the Milky Way’s life, bringing with it, some $\sim 2\times 10^{9}M_{\odot}$ of stellar mass. The current consensus on the accreted nature of the Canis Major is rather sceptical given the disc origin of the various “streams” associated with it (see e.g. Xu et al., 2015; Price-Whelan et al., 2015; de Boer et al., 2018; Sheffield et al., 2018; Laporte et al., 2020). Most importantly, however, neither of the two groups of Galactic GCs identified by Kruijssen et al. (2019b) exhibits coherent orbital behaviour and thus both are instead most likely amalgamations of parts of different mergers. The analysis of Kruijssen et al. (2019b) is trail-blazing but falls short due to the limited quality of the age-metallicity information for the Galactic GCs and the neglect of the orbital information. An attempt to rectify the latter is made in Kruijssen et al. (2020), which re-uses the IOM classification made by Massari et al. (2019). In a similar vein, Forbes (2020) starts with GC groups constructed by Massari et al. (2019) and tweaks them very slightly using the age-metallicity information at hand.
Figure 14: Groups of GCs in action-angle space inferred from chemodynamical modelling. The coloured symbols are the observed GCs: different colours indicate different groups, with each GC coloured by its most likely group. The accreted groups of GCs are indicated with their $1-,2-,3-\sigma$ confidence intervals. [Reproduced from Callingham et al. (2022)].
@@
Compared to all preceding attempts, the study of Callingham et al. (2022) stands out as the only unbiased GC classification scheme. Instead of hand-picking groups of clusters in the chosen parameter space, they devise a method similar to Gaussian Mixture Modelling. For the Milky Way GCs, ages, metallicities, energy, and three orbital actions are modeled with component mixtures to obtain objective membership probabilities for each GC (see Fig. 14). This innovative approach is demonstrated to work well for simulated data. Unfortunately, Callingham et al. (2022) also reports a large degree of overlap between GCs across all dimensions studied. To mitigate against the arising degeneracies, Callingham et al. (2022) decide to initialize their mixture models to the GC classes reported by Massari et al. (2019).
Concerned with the separation between the GCs born in-situ and those formed in dwarfs and later accreted, Belokurov and Kravtsov (2023) go back to the idea of the “critical energy” highlighted in Myeong et al. (2018c). They show that in the stellar halo too, a rather sharp boundary exists in $E,L_{z}$ space which divides stars into mostly accreted and mostly in-situ. This division is mapped by the clear change in the overall levels of [Al/Fe] ratio amongst the field stars. Using the $E,L_{z}$ boundary delineated by the [Al/Fe] levels, Belokurov and Kravtsov (2024) classify all Milky Way GCs into two classes and find 106 out of 164 to be of in-situ origin, and 58 to be accreted. Therefore the most up-to-date GS/E membership of $\sim 20$ GCs translates into $\sim 1/3$ of the total accreted cluster population in the Galaxy. Note that although the proposed GC classification scheme based on their position in the $E,L_{z}$ plane has recently been independently verified in Chen and Gnedin (2024), Belokurov and Kravtsov (2024) caution that some $\approx 10\%$ of the objects classified as in-situ could be of accreted origin. This will bring down the GS/E fractional contribution to the accreted GC population in the MW.
Finally, Valenzuela et al. (2023) recently argue that the GS/E GCs themselves have a bimodal age distribution, which they suggest is the result of a “wet” GS/E merger that both brings in its own GCs while also forming new GCs in the process of the gas-rich merger with the Milky Way. This intriguing scenario allows for the gas mass of the GS/E progenitor to be estimated based on the properties of its GCs.
@@ 3.5 Simulations
While the above reflects a more tailored approach to the simulations of a Milky Way-GS/E merger, other works have focused on large suites of Milky Way-mass galaxies simulated in a cosmological context. In this case, an ‘ideal’ Milky Way-GS/E system is harder to find, but such analyses are essential in order to put the Milky Way assembly history in context with the wider galaxy population. Fattahi et al. (2019) use the Auriga suite of simulations to show that approximately one-third of the (28) haloes have a dominant stellar halo component similar to the observed Milky Way. From this subset, Fattahi et al. (2019) deduce that these merger events typically occur 6-10 Gyr ago, with merging fragments of stellar mass $M_{\rm star}\sim 10^{9}-10^{10}M_{\odot}$. Orkney et al. (2023) extend the analysis of the GS/E events in Auriga and highlight the diversity of possible progenitor systems, with a broad range of stellar, dark matter masses and rotational properties. Mackereth et al. (2019b) use the EAGLE simulation suite to show that merger events with a similar pattern in chemistry and dynamics to the real GS/E are typically related to massive dwarfs with $M_{\rm star}\sim 10^{8.5}-10^{9}M_{\odot}$. Moreover, they argue that the high eccentricity of the GS/E stars today suggests the merger would have taken place at $z\lesssim 1.5$. One particular Milky Way analogue within the EAGLE suite is additionally studied in greater detail by Bignone et al. (2019). Elias et al. (2020) show that many of the conclusions reached through studies of GS/E-like events in suites of zoom-in simulations also hold in a much larger but lower-resolution sample from Illustris. Finally, Dillamore et al. (2022) uses the ARTEMIS simulations to again analyse GS/E-like events in cosmological simulations. In agreement with Fattahi et al. (2019) they find that typically a third of the Milky Way-mass systems contain a feature similar to the GS/E.
@@
Moving forward, a combined effort of tailored Milky Way-GS/E mergers (e.g. Koppelman et al., 2020; Naidu et al., 2021) and analyses of cosmological simulation suites (e.g. Fattahi et al., 2019; Mackereth et al., 2019b; Dillamore et al., 2022) are needed to (a) link the observed properties of the GS/E to its initial state and co-evolution with the Milky Way, and (b) calibrate our various observational techniques to infer these properties. An important step in the right direction are the works of Amarante et al. (2022) and Rey et al. (2023) who create hydro-dynamical models of the GS/E-Milky Way encounter. Amarante et al. (2022) build tailor-made single merger models with realistic star-formation recipes and detailed chemistry. They show that many previously detected halo substructures can indeed be attributed to the single GS/E-like event. Rey et al. (2023) pioneers a new powerful and flexible way to carry out inference with Cosmological hydro-dynamical simulations. They use genetic modifications of the simulation’s initial conditions (see Roth et al., 2016; Rey and Pontzen, 2018) to modify the properties (such as the mass ratio) of the GS/E merger. Rey et al. (2023) take a Milky Way zoom-in simulation from the VINTERGATAN suite and show how the properties of both the host and the resulting debris cloud change as the merger strength is varied. The Milky Way suffers more dramatic truncation and becomes more bulge-dominated as the GS/E mass is increased (see Fig. 15). In the absence of detailed chemistry, Rey et al. (2023) find a pronounced kinematic degeneracy for the stellar halo: different mergers and different mixtures of accreted and in-situ stars end up producing similar looking radially-biased “sausage”-like structures.
It is worth emphasizing a sobering point regarding simulated ‘Milky Way-like’ galaxies. While they can be extremely helpful for understanding and interpreting the observational data, they are not, and never will be, the ‘real’ Milky Way. Moreover, the selection effects present in the data are not often (and rarely fully) accounted for in the simulations. Thus, great care must be taken when using the simulations to draw conclusions about our own Galaxy.
@@ 3.6.2 The mass budget of the GS/E progenitor
The most recent stellar mass estimates of the GS/E are typically a few ×$\times 10^{8}M_{\odot}$ (e.g. Mackereth and Bovy, 2020; Lane et al., 2023), which is almost an order of magnitude lower than earlier estimates of $\sim 10^{9}M_{\odot}$ (e.g. Belokurov et al., 2018b; Helmi et al., 2018; Feuillet et al., 2020). This is likely due to a greater purity of GS/E stars in more recent studies, but, even now, a robust stellar mass estimate is lacking. Nonetheless, a lower mass estimate for GS/E has several important consequences. For example, it would imply less damage to the pre-existing disc — this would agree with observations that the high-$\alpha$ disc is still intact. However, simulated analogues of the Milky Way-GS/E encounter would need to be updated in light of this seemingly preferred lower GS/E mass.
@@
Finally, it is also worth considering what the gas content of the GS/E progenitor was. The early nature of the merger strongly suggests that the progenitor was gas-rich, but where and when was the gas lost? Could it contribute to (or re-start) the Galactic star formation? Does the two infall model of the Galaxy (Chiappini et al., 1997) still hold weight in light of the apparent gas-rich GS/E merger? Can the GS/E GC population be used to estimate the gas mass of the progenitor (Valenzuela et al., 2023)?
@@ 3.6.3 Initial state of the progenitor
Although much work is needed to characterize the GS/E debris today, ultimately we also want to know what the initial state of the progenitor was (i.e. before any interactions with the Milky Way). How much information is encoded in the $z=0$ debris? For example, if the progenitor was a disc galaxy with a large satellite (Sequoia?), broadly similar to the LMC and SMC, it could produce a variety of halo substructures with distinct properties. However, even without a complex morphology and kinematics, high-mass events tend to break up into a large number of clumps of the $E,L_{z}$ space (see e.g. Jean-Baptiste et al., 2017; Belokurov et al., 2023). Indeed, the ‘chevrons’ in $r-V_{r}$ space uncovered by Belokurov et al. (2023) could result from the phase-mixing of the massive GS/E. However, Dillamore et al. (2023a) recently showed that such stellar substructures in the local halo could also generated by bar resonances.
@@ 3.6.4 Debris distribution in the Galactic halo
Most of the analysis of the GS/E debris has been limited to the inner ($5\lesssim r/\mathrm{kpc}\lesssim 30$) halo. This, in part, is because here we have the richest multi-dimensional data, but also because the high eccentricity of the GS/E orbit implies there are plenty of GS/E stars at small Galactic radii. However, there is not much evidence (so far) for GS/E stars at low total $E$. How deep a satellite can sink in the halo depends not only on its mass, but also on (i) the density of the host and the satellite, and (ii) the initial amount of angular momentum (Amorisco, 2017; Vasiliev et al., 2022). The low $E$ regime is complicated by the overwhelming in-situ population, but there is potentially a rich amount of uncharted information about the GS/E lurking in the Milky Way’s depths.
At the other ‘end’ of the Galaxy, there is still much to learn about the GS/E stars at large distances. Here, shells and earlier stripped GS/E stars are likely prevalent in the halo. Indeed, the distant ‘echoes’ of the GS/E are now starting to be discovered (Chandra et al., 2023), and with increasing amounts of data at large Galactic radii, this will likely be a fruitful expedition in future studies. Mapping the GS/E debris over a wide range of radii (or energies) will be vital in order to reconstruct the full dynamical history of the Milky Way-GS/E interaction.
@@ 3.6.5 A complete re-interpretation
Although the vast majority of the works described above advocate (or assume?) that the GS/E is a single, early, and massive progenitor, this is not the only interpretation in the literature. In particular, in a series of recent papers (Donlon et al., 2022; Donlon and Newberg, 2023; Donlon et al., 2024), Donlon et al. argue that the GS/E in fact comprises debris from several lower mass radial mergers, which moreover, were accreted fairly recently. This alternative view, coupled with the large uncertainty in several of the GS/E properties (e.g. total stellar mass, time of accretion) shows that we are still far from a robust characterization of this stellar halo component. Perhaps the best way to avoid the degeneracy between total mass and the number of progenitors is the use of the well-known stellar mass-metallicity (MZR) relation (as several lower mass fragments will have lower metallicity than a single higher mass system). However, the biases associated with different selection methods (see e.g. Carrillo et al., 2024) coupled with uncertainties in the time evolution of the MZR has hampered this crucial line of evidence (see also Section 5.3). Nonetheless, it is clear that a scenario accounting for all of the chemodynamical evidence is required in order to fully solve the GS/E puzzle.
@@ 4.1 In-situ stellar halo
Figure 16: Left: Azimuthal velocity ($V_{\phi}$) as a function of age for local metal-rich stars selected from Gaia DR2. Middle: Distribution of ages for stars in different bins of $V_{\phi}$ and [Fe/H]. Disc stars (red) are metal-rich with high $V_{\phi}$, Splash (black) is metal-rich with low $V_{\phi}$, and GS/E (blue) stars are metal-poor(er) with low $V_{\phi}$. Right: Cumulative age distribution. Belokurov et al. (2020) argue that the truncated age distribution of the Splash stars (indicated by a vertical thick red line) can put a strong constraint on the time of the last major merger (if indeed the Splash is caused by this merger event)
. [Adapted from Belokurov et al. 2020]